OpenStudy (fatimaj13):

cos theta -cos theta sin^2 theta. I know the answer is cos ^3 theta but i'm not sure how i got the answer

OpenStudy (holsteremission):

\[\cos\theta-\cos\theta\sin^2\theta=\cos\theta\left(1-\sin^2\theta\right)\]If it's unclear what happened there, think of the left hand side as \[a-ab^2=a(1-b^2)\] Now, what do you know about \(1-\sin^2\theta\)?

OpenStudy (welshfella):

1 - sin^2 theta = cos^2 theta

OpenStudy (welshfella):

cos cos theta ( 1 - sin^2 theta) = cos^3 theta

OpenStudy (welshfella):

- ignore the first cos!!